OpenStudy (anonymous):

Please help Suppose the force between the Earth and Moon were electrical instead of gravitational, with the Earth having a positive charge and the Moon having a negative one. If the magnitude of each charge were proportional to the respective body’s mass, find the Q/m ratio required for the moon to follow its present orbit of 3.84 x 108 m radius with its period of 27.3 days. The Earth’s mass is 5.98 x 1024 kg, and the moon’s mass is 7.3 x 1022 kg.

OpenStudy (anonymous):

Calculate the centripetal force required first.

OpenStudy (anonymous):

would I make it equal to the electric force after?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

would I have to do anything after that?

OpenStudy (anonymous):

also would the mass of earth and moon be the charges on them?

OpenStudy (anonymous):

No. You'll have \[Q_{moon} = C \cdot M_{moon} \] and \[Q_{earth} = C \cdot M_{earth} \] so \[Q_{moon} Q_{earth} = C^2 M_{moon} M_{earth} \] the question wants you to solve for C.

OpenStudy (anonymous):

so the charge of the moon is equal to the centripetal force multiplied by the mass of the moon

OpenStudy (anonymous):

No, of course not. The C is just a constant.

OpenStudy (anonymous):

so I will isolate for c then

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so I will isolate c first, the find centripetal force then electric

OpenStudy (anonymous):

I'm not sure what you mean, but just try it.

OpenStudy (anonymous):

so when I find C, the constant, will that be the answer? because I don't know were finding centripetal force and electrical force come in

OpenStudy (anonymous):

Yes it will. You can find the centripetal force and set that equal to the electric force. That's one equation, and the unknown is C.

OpenStudy (anonymous):

so the equation will be Me Mm 4pirf^2 = (kQ1Q2/ r^2) C

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

What is the expression for centripetal force?

OpenStudy (anonymous):

v^2/r = 4pi^2 rf^2=4pi^2r/T^2

OpenStudy (anonymous):

That's centripetal acceleration. You need to multiply by the mass of the object that's moving in the circle.

OpenStudy (anonymous):

I forgot the mass in each equation tho.

OpenStudy (anonymous):

so it will be Mm 4pirf^2 = (kQ1Q2/ r^2) C , but Mm will be only the mass of the moon

OpenStudy (anonymous):

The left side is right. The right side has no C in it. The C comes in when you replace the Q's with M's, like I did above.

OpenStudy (anonymous):

so with that equation I will solve for Q for the earth and moon

OpenStudy (anonymous):

Just read what I wrote above. Qe Qm = C^2 Me Mm

OpenStudy (anonymous):

ok, I will try it then then come back here to see if I'm doing it right and if I get what you mean

OpenStudy (anonymous):

so i found Q for the moon, but for the earth i don't have r or f so i don't know how to do it. i used the equation Q= m4pi^2r^3f^2/k

OpenStudy (anonymous):

\[F_c = 4\pi^2M_{moon}fR = k \frac{Q_{earth}Q_{moon} }{R^2}= C^2k \frac{M_{earth}M_{moon}}{R^2}\] Can you do it from here?

OpenStudy (anonymous):

yes. sorry

OpenStudy (anonymous):

Quite alright.

OpenStudy (anonymous):

ugh I'm very sorry. But what is messing me up is the Q and M. Because I don't know what I have to sub in for Q, and solve for. I know I have to solve for C for the last equation, but what about the other two, am I solving for Fe for the other two and making it equal to the one with c and isolating c

OpenStudy (anonymous):

You are setting the first equal to the last. You know all of those values except for C.

OpenStudy (anonymous):

Well, the second -- the one that starts with 4 pi ^2